# HALT: Hallucination Assessment via Latent Testing

### v1

@@ 5 Ablation Studies
To understand the design and limitations underlying our hallucination detector, we conduct several ablation studies analyzing how different component affect performance. In particular, we explore (1) which transformer layers provide the most informative representations for hallucination detection, (2) how model scale influences both classification separability and ranking quality under selective deployment, (3) how different detector architectures and representation aggregation strategies impact performance, and (4) whether we can use a detector trained on the question and answer to outperform a detector trained on only the question, during generation for parallel evaluation. Unless otherwise stated, all ablations are performed using the Qwen family of models.
@@ 5.1 Effect of Intermediate Layers on Performance
Figure 4: AUROC and AURAC across layers 0–28 for Qwen2.5-7B on Webquestions. Both metrics peak around layer 20, suggesting that intermediate-to-late layers encode the most informative signals for uncertainty detection.
Recent work has shown that different layers in transformers encode qualitatively different information: early layers capture surface-level features, while deeper layers encode higher-level semantics and factual knowledge (Mir, 2025; Ni et al., 2025). We hypothesize that intermediate layers, which balance low-level token representations with high-level semantic understanding, may provide the strongest signals for hallucination detection. To test this, we train separate detectors on hidden states extracted from each layer and compare their performance.
Figure 4 shows AUROC and AURAC across all 29 layers of Qwen2.5-7B. Both metrics increase monotonically from early layers and peak around layer 20, achieving AUROC of 0.877 and AURAC of 0.643. Performance then gradually declines in the final layers. This pattern suggests that intermediate-to-late layers, which are believed to encode higher-level semantic and factual information, provide the strongest signals for detecting hallucinations. We observe similar trends for LLaMA-2-7B and Gemma-3-4B (see Appendix C in the Appendix).
5.2 Effect of Model Size on Ranking Quality
Figure 5: Effect of model scale (7B, 14B, 32B) on AUROC and AURAC. While AUROC remains largely stable across scales, AURAC increases consistently, indicating improved ranking quality under selective deployment for larger models.
We study the effect of model scale on selective deployment performance by evaluating Qwen2.5 models at 7B, 14B, and 32B parameters on the same question sets. As shown in Figure 5, AUROC remains largely stable across scales on both WebQuestions and MMLU-Pro, indicating that overall classification separability changes little with model size. In contrast, AURAC improves consistently with scale, reflecting substantially better ranking quality for selective prediction. The improvement is most pronounced on MMLU-Pro, where the 32B model achieves a 12.7-point AURAC gain over the 7B model. This trend indicates that larger models produce confidence scores that are better aligned with correctness, enabling more effective filtering of unreliable responses under selective deployment.
@@ 5.3 Effect of Detector Architecture
Figure 6: Comparison of detector architectures on TriviaQA (Qwen2.5-7B, layer 19). Among MLP variants, last-token pooling achieves the highest AUROC (93.2%), outperforming mean pooling (92.1%) and max pooling (90.9%). The transformer architecture achieves the best AUROC (93.95%) with attention pooling.
We compare the two detector architectures described in Section 3: an MLP with input-stage pooling and a lightweight transformer with attention pooling (Table 2). For the MLP, we further compare three pooling strategies: last token, mean pooling, and max pooling. All experiments use hidden states from layer 19 of Qwen2.5-7B on TriviaQA.
As shown in Figure 6, last-token pooling consistently outperforms mean and max pooling across all metrics, achieving 93.2% AUROC compared to 92.1% and 90.9%, respectively. The transformer architecture with attention pooling achieves the highest AUROC (93.95%), slightly outperforming the best MLP variant. Based on these results, we use the transformer architecture with attention pooling in our main experiments.
5.4 Effect of Number of Answer Tokens Used
Figure 7: Effect of answer token inclusion on detection performance. We vary the percentage of answer tokens used (5%–100%) alongside all question tokens. Both AUROC and AURAC increase monotonically with more answer tokens.
Our main experiments compare two settings: using only question tokens versus using question plus all answer tokens. This ablation fills the gap between these two extremes by progressively increasing the proportion of answer tokens. Specifically, we train detectors using question tokens plus the first $X\%$ of answer tokens, where $X$ ranges from 5% to 100%. Figure 7 shows the results on TriviaQA using Qwen-2.5-7B.
Both AUROC and AURAC increase monotonically with the token threshold. At 5% answer tokens, the detector achieves 82.2% AUROC; at 100%, this rises to 87.1%. AURAC follows a similar trend, increasing from 0.628 to 0.655. These results confirm that answer tokens provide additional signal for hallucination detection, and this signal accumulates as more tokens are included.
@@ 6 Discussion
Our experiments focus on single-turn question-answering tasks with relatively short contexts. In multi-turn dialogues or long-context scenarios, the intermediate representations of the current question may be influenced by the preceding conversation history, potentially degrading detector performance. Adapting our method to such settings may require collecting training data that reflects these longer contexts, and employing compression or sliding window strategies to select informative residual representations for training and inference. However, we emphasize that the length of the generated answer does not affect our method, since we are able to achieve superior results in most settings depending only on question token representations, and can make predictions before answer tokens are produced.
@@
In Section 5.2 (Figure 5), we find that our method performs better when applied to larger models. A similar phenomenon has been reported in prior work (Du et al., 2024; Kadavath et al., 2022). We therefore conjecture that our method generalizes to stronger, larger LLMs. One possible explanation is that larger models have greater capacity and may perform richer latent computation beyond next-token prediction. As a result, the final text can be less faithful to the intermediate state. This can increase the gap between intermediate representations and surface outputs, making intermediate features more informative for our detector thereby improving its effectiveness on larger models.
@@
Beyond hallucination detection and selective prediction, our method may benefit reinforcement learning for LLMs. In tree search algorithms such as MCTS, the uncertainty signal from our detector could guide exploration: branches where the model is uncertain warrant more exploration, while confident branches can be pruned early. This could help prevent the model from converging to local minima driven by overconfident but incorrect predictions, leading to more robust policy optimization.
@@
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix B Dataset Statistics
Table 5: Correctness label statistics across models and benchmarks. Each model has two rows: the first shows the number of correct/incorrect samples, and the second shows the corresponding accuracy (%). For NQ-Open and TriviaQA, 20K samples were randomly selected from the original training set; the remaining benchmarks use full training sets. All benchmarks include the complete test set. The counts shown represent the combined training subset and test set.
@@
Appendix C Ablation Study 2: Layer Ablation on Additional Models
We extend our layer ablation study to LLaMA-2-7B and Gemma-3-4B to verify that the trend observed in Qwen2.5-7B generalizes across model families. Figure 8 shows AUROC and AURAC across all layers for both models on the WebQuestions dataset.
For LLaMA-2-7B (32 layers), performance peaks around layer 14, with AUROC and AURAC declining in both earlier and later layers. For Gemma-3-4B (26 layers), the peak occurs around layer 18–20. Interestingly, the Gemma-3-4B curve exhibits notable fluctuations across layers, suggesting that different layers may encode qualitatively different information; this warrants further investigation. In both cases, intermediate-to-late layers outperform early layers and the final layers, confirming that the optimal layer for hallucination detection is not the last layer but rather an intermediate one. This pattern aligns with findings in interpretability literature suggesting that middle layers encode richer semantic representations.
@@ Appendix C Ablation Study 2: Layer Ablation on Additional Models
Figure 8: Layer ablation results for LLaMA-2-7B and Gemma-3-4B on WebQuestions. Both models show peak performance at intermediate layers, consistent with the trend observed in Qwen2.5-7B.

### v2

@@ 5 Ablation Studies
To understand the design choices underlying our hallucination detector, we conduct several ablation studies. In particular, we explore: (1) which transformer layers provide the most informative representations for hallucination detection, (2) how model scale influences classification and ranking quality, (3) how different detector architectures and pooling strategies impact performance, and (4) how detection performance varies with the number of answer tokens available at inference time. Unless otherwise stated, all ablations are performed using Qwen2.5-7B-Instruct.
@@ 5.1 Effect of Intermediate Layers on Performance
Figure 4: AUROC and AURAC across layers 0–28 for Qwen2.5-7B on WebQuestions. Both metrics peak around layer 20.
As discussed in Section 3, we hypothesize that intermediate layers are more informative than the final layer for hallucination detection. To validate this and identify the optimal layer, we train separate detectors on hidden states extracted from each layer and compare their performance. This design is motivated by prior work showing that different transformer layers encode qualitatively different information (Mir, 2025; Ni et al., 2025).
Figure 4 shows AUROC and AURAC across layers 0–28 of Qwen2.5-7B. Both metrics show an overall upward trend with minor fluctuations, peaking around layer 20 (AUROC 0.877, AURAC 0.643), before gradually declining in later layers. This pattern suggests that intermediate-to-late layers provide the strongest signals for detecting hallucinations. We attribute the decline in final layers to the projection toward token space, which discards internal confidence signals that are unnecessary for next-token prediction. We observe similar trends for LLaMA-2-7B and Gemma-3-4B (see Appendix C).
5.2 Effect of Model Scale
Figure 5: Effect of model scale (7B, 14B, 32B) on AUROC and AURAC. AUROC remains stable on WebQuestions and improves on MMLU-Pro. AURAC improves consistently on both datasets.
In Section 1, we hypothesized that larger models exhibit a greater gap between internal representations and surface outputs, making our detector more effective. To verify this, we evaluate Qwen2.5 models at 7B, 14B, and 32B on WebQuestions and MMLU-Pro. As shown in Figure 5, AUROC remains stable on WebQuestions and improves on MMLU-Pro. AURAC improves consistently, with the 32B model achieving a 12.7-point gain over 7B on MMLU-Pro. Both trends confirm that larger models yield stronger detection signals and better-calibrated confidence scores, supporting our hypothesis. Given these scaling trends, we expect our method to be even more effective on frontier models with hundreds of billions of parameters.
@@ 5.3 Effect of Detector Architecture
Figure 6: Comparison of detector architectures on TriviaQA (Qwen2.5-7B, layer 19). Last-token pooling performs best among MLP variants; the transformer architecture with attention pooling achieves the highest overall performance.
In Section 3, we described two candidate architectures for the detector: an MLP with input-stage pooling and a lightweight transformer with attention pooling (Table 2). For the MLP, multiple pooling strategies are possible: last token, mean pooling, and max pooling. Here we compare these design choices to identify the best configuration. All experiments use hidden states from layer 19 of Qwen2.5-7B on TriviaQA.
As shown in Figure 6, among MLP variants, last-token pooling performs best, with the largest gain in Accuracy (3.6 points over max pooling) and improvements of 1.1–2.3 points in AUROC and 0.8–3.6 points in AURAC. The transformer architecture further improves upon this, achieving comparable Accuracy while gaining 0.8 points in AUROC and 0.6 points in AURAC. Based on these results, we adopt the transformer architecture in our main experiments.
5.4 Effect of Answer Token Availability
Figure 7: Effect of answer token availability on detection performance (Qwen2.5-7B, MMLU-Pro). Both AUROC and AURAC increase monotonically as more answer tokens are included.
Our main experiments compare question-only and question-plus-answer settings. Here we explore an interesting setting: what if we evaluate the detector during generation, using only the answer tokens produced so far? We train a detector on question plus all answer tokens, then evaluate using only the first $x\%$, where $x$ ranges from 5% to 100%. We use MMLU-Pro for this ablation because its answers involve multi-step reasoning, unlike TriviaQA and NQ-Open, where answers are often single tokens.
As shown in Figure 7, both AUROC and AURAC increase monotonically with the token threshold, exhibiting highly consistent behavior across the entire range. AUROC improves by 4.9 points and AURAC by 2.8 points from 5% to 100% tokens. This confirms that answer tokens carry cumulative signals for hallucination detection. This enables flexible deployment. Beyond question-only detection that runs fully in parallel with generation, practitioners can trigger detection partway through the response, achieving higher accuracy without incurring additional latency, which is critical for real-time user experience.
@@ 6 Discussion
Correctness as a Proxy for Hallucination
We use correctness as a proxy for hallucination, which is standard practice in this field but warrants discussion. Strictly speaking, our method detects cases where the model is uncertain yet still produces an output, rather than all types of hallucinations. Models can produce errors for multiple reasons. First, a lack of relevant knowledge, which our method can detect. Second, systematic errors caused by incorrect training data, which our method cannot detect. Third, logical errors during reasoning, which our method cannot detect. Conflating these distinct mechanisms is not appropriate. Furthermore, for ambiguous questions, underspecified queries, or questions with multiple valid answers, an incorrect label does not necessarily indicate hallucination. Our datasets (TriviaQA, NQ-Open, MMLU-Pro, WebQuestions) primarily consist of factoid QA with well-defined answers, where this limitation has less impact. Therefore, our reported results can be viewed as a lower bound on detection capability. The fact that we achieve strong performance despite correctness, including error types that our method is not designed to detect, suggests that model uncertainty is a major source of LLM errors.
Data Imbalance
Some model-dataset combinations exhibit severe label imbalance, as indicated by * in Table 4. For example, LLaMA-2 achieves only 15.4% accuracy on MMLU-Pro, meaning that correct answers constitute a small minority of the training data. Although AUROC is more robust to class imbalance than accuracy (since it measures ranking quality rather than classification at a fixed threshold), results under extreme imbalance should still be interpreted with caution. Notably, on settings with more balanced labels (such as Qwen across all datasets), our method achieves the strongest results across all benchmarks, suggesting that the method itself is effective and the weaker performance on imbalanced settings is attributable to data limitations rather than methodological flaws.
@@
Our experiments focus on single-turn, closed-book question-answering tasks. We do not test on multi-turn dialogues, tool use, code generation, or summarization. We choose QA because it is a standard benchmark for hallucination detection research with clear ground-truth labels for evaluation. Adapting our method to other scenarios requires collecting corresponding training data; in long-context settings, compression or sliding-window strategies may be needed to select informative representations. The core assumption of our method (that intermediate layers encode uncertainty signals) is independent of task type, and we expect it to generalize to other generation tasks, though this requires further validation.
@@
In Section 5.2 (Figure 5), we find that our method performs better when applied to larger models. A similar phenomenon has been reported in prior work (Du et al., 2024; Kadavath et al., 2022). We therefore conjecture that our method generalizes to stronger, larger LLMs. One possible explanation is that larger models have greater capacity and may perform richer latent computation beyond next-token prediction. As a result, the final text can be less faithful to the intermediate state. This can increase the gap between intermediate representations and surface outputs, making intermediate features more informative for our detector, thereby improving its effectiveness on larger models.
@@
Beyond hallucination detection and selective prediction, our method may benefit reinforcement learning for LLMs. In tree search algorithms such as Monte Carlo tree search (MCTS), the uncertainty signal from our detector could guide exploration: branches where the model is uncertain warrant more exploration, while confident branches can be pruned early. This could help prevent the model from converging to local minima driven by overconfident but incorrect predictions, leading to more stable policy optimization.
@@
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Acknowledgments
HY was partially supported by the US National Science Foundation under awards IIS-2520978, GEO/RISE-5239902, the Office of Naval Research Award N00014-23-1-2007, DOE (ASCR) Award DE-SC0026052, and the DARPA D24AP00325-00. Approved for public release; distribution is unlimited.
@@ Appendix B Dataset Statistics
Table 5: Correctness label statistics across models and benchmarks. Accuracy (%) is shown for each model-dataset pair.
@@
Appendix C Layer Ablation on Additional Models
We extend the layer ablation study from Section 5.1 to LLaMA-2-7B and Gemma-3-4B to verify that the trend observed in Qwen2.5-7B generalizes across model families. Figure 8 shows AUROC and AURAC across all layers for both models on the WebQuestions dataset.
For LLaMA-2-7B (32 layers), performance peaks around layer 14, with AUROC and AURAC declining in both earlier and later layers. For Gemma-3-4B (35 layers), the peak occurs around layer 28. Interestingly, the Gemma-3-4B curve exhibits notable fluctuations across layers, suggesting that different layers may encode qualitatively different information; this warrants further investigation. In both cases, intermediate-to-late layers outperform early layers and the final layers, confirming that the optimal layer for hallucination detection is not the last layer but rather an intermediate one. This pattern aligns with findings in interpretability literature suggesting that middle layers encode richer semantic representations.
@@ Appendix C Layer Ablation on Additional Models
Figure 8: Layer ablation results for LLaMA-2-7B and Gemma-3-4B on WebQuestions. Peak performance occurs at intermediate layers.